How many cups of grapefruit juice must be added to 40 cups of a punch that is 2% grapefruit juice to obtain a punch that is 10% grapefruit juice?
Answer by oberobic(2304)

To solve mixture problems, you have to determine how much "pure" stuff you have and how much you need in the final mix.
.
You have 40 cups of 2% grapefruit juice.
40*.02 = .8 cup of pure grapefruit juice, or 8/10 of a cup.
.
You have an unlimited supply of 100% grapefruit juice that you have to mix with the 2% to obtain a final amount of juice mixture that is 10% grapefruit juice.
.
Looking at this logically, if you had 10% juice, you would have to have 4 cups of pure juice in the 40 cups of punch. But, everything you add increases the volume, so you have to model it.
.
x = amount of pure juice to add
40+x = amount of 10% punch you will have in the end.
.
Pure grapefruit juice in the final produce = 10%*(40+x).
.
Your ingredients are:
2%*40 + 100%*x = 10%*(40+x)
.02*40 + x = .1(40+x)
.
Multiply by 100 to remove percents
.
2*40 + 100x = 10(40+x)
80 + 100x = 400 + 10x
90x = 320
x = 320/90
x = 32/9 cups
x = 3 5/9 cups to add to the 40 cups you have
.
Check your answer by determining how much pure stuff you have vs. how much you ought to have.
.
You have 43 5/9 cups of total mix.
= 43.555 cups
.
You have 8/10 + 32/9 cups of pure stuff.
= 4.3555 cups
.
So, the final mixture is indeed 10% grapefruit juice.
.
Answer:
Add 3 5/9 cups of pure grapefruit juice.